### v2

@@ Abstract
%Fulfilling privacy requirements is essential for software development companies and is largely motivated by customer demands and high fines for non-compliance. However, past research showed that this is a challenging endeavor: While software developers are held responsible for implementing privacy requirements, they often lack the critical background knowledge and support to fulfill this task. To shed light on this issue, we conducted a lab programming study with 30 developers that lasted an average of 3 hours and 23 minutes. Participants were divided into three groups: a control group, a privacy prompted group, and an expert-support group with the option of contacting a privacy expert with questions. Participants completed four programming tasks in which they had to decide whether and how to ensure privacy-compliant solutions, followed by an interview. %We found that although most developers exhibited minimal confidence in their solutions, interestingly, they did not seek online assistance or contact the study’s privacy expert. Participants often referenced security, suggesting privacy was a secondary concern behind functionality and security, even when specifically prompted for privacy.
While protecting user data is essential, software developers often fail to fulfill privacy requirements. However, the reasons why they struggle with privacy-compliant implementation remain unclear. Is it due to a lack of knowledge, or is it because of insufficient support? To provide foundational insights in this field, we conducted a qualitative 5-hour programming study with 30 professional software developers implementing 3 privacy-sensitive programming tasks that were designed with GDPR compliance in mind. To explore if and how developers implement privacy requirements, participants were divided into 3 groups: control, privacy prompted, and privacy expert-supported. After task completion, we conducted follow-up interviews. % Alarmingly, almost all participants submitted non-GDPR-compliant solutions (79/90). In particular, none of the 3 tasks were solved privacy-compliant by all 30 participants, with the non-prompted group having the lowest number of 3 out of 30 privacy-compliant solution attempts. Privacy prompting and expert support only slightly improved participants’ submissions, with 6/30 and 8/30 privacy-compliant attempts, respectively. In fact, all participants reported severe issues addressing common privacy requirements such as purpose limitation, user consent, or data minimization. Counterintuitively, although most developers exhibited minimal confidence in their solutions, they rarely sought online assistance or contacted the privacy expert, with only 4 out of 10 expert-supported participants explicitly asking for compliance confirmation. Instead, participants often relied on existing implementations and focused on implementing functionality and security first. %Instead, they often referenced security, suggesting privacy was a secondary concern.
%%The abstract is a short summary of the work to be presented in the %article.
@@ 1 Introduction
Several regulations have been passed to reign in tech companies’ data collection practices and empower users to regain control over their personal data. The best-known ones are Japan’s Act on the Protection of Personal Information (APPI) from 2003, South Africa’s Protection of Personal Information Act (POPI Act) from 2013, the General Data Protection Regulation (GDPR) [1] (2018) by the European Union, and the California Consumer Privacy Act CCPA [2] (2018) by the state of California. Consequently, user privacy has become an essential requirement that software development companies must accomplish. Severe fines [3], fear of bad publicity [4], and customer expectations [5] all lead to an increased focus on developing privacy-preserving software products. Still, since the introduction of e.g., the GDPR in 2018, more than 5 billion euros have been issued for violations [6].
While larger companies often have access to legal experts to support software developers, their communication with software development teams tasked with fulfilling these requirements is often limited, suggesting a “communication gap” [5]. Although smaller companies are hit harder by fines for privacy non-compliance, they are often unable to hire privacy experts due to limited budget restrictions. As a result, fulfilling privacy requirements has become a task handled solely by the software development teams.
However, past research showed that this has not (yet) translated into the actual practices of individual developers (e.g., [3, 7, 8, 9, 10, 11, 12]), raising the question: are developers unaware of privacy requirements, or is the task of fulfilling them lacking support?
To address this research gap, we conducted a lab programming study with $30$ professional software developers implementing requirements for a health application, lasting 3 hours and 23 minutes on average. We asked participants to complete 4 programming tasks, with 3 of them requiring a GDPR-compliant implementation: (i) a warm-up task consisting of fixing an application’s backup function, (ii) processing a deletion request from one of the application’s users, (iii) implementing a search function for doctor users of the application, and lastly, (iv) implementing an e-mail function to send user advertisements and obtain user consent for a privacy-compliant solution. With this work, we explored the following research questions:
@@
To explore developers’ privacy awareness and address the communication gap [5], we separated the participants into three groups: (i) a control group without mentioning privacy at all, (ii) a group prompted on privacy-compliant implementation, and (iii) a chatting group having access to a privacy expert. To further understand participants’ implementation challenges with privacy requirements during the study, we conducted follow-up interviews.
With this work, we contribute foundational insights into a highly complex problem.
Our findings suggest a lack of privacy awareness, the perception that privacy requirements are overly complex to implement, and a reluctance to consult privacy experts, as developers often fear being a burden or inconveniencing them. Participants cited a lack of knowledge and experience and relied on preexisting implementations. Still, most interactions with experts were limited to confirming compliance rather than seeking guidance on privacy requirements.
@@ 2.1 Regulations
Previous research [13, 14, 3, 15, 16, 17] provided comprehensive insights into how GDPR and similar regulations shape the software development process. Marjanov et al. [14] emphasized the necessity of both technical and organizational measures to meet legal standards, identifying specific danger points and conditions that lead to data breaches. This data-driven approach offered developers crucial insights into compliance risks and mitigation strategies. Saemann et al. [3] further explored the impact of regulations through the analysis of GDPR fines, highlighting common compliance failures, particularly during data collection. They advocated for the integration of Privacy-by-Design principles early in the development cycle to enhance data protection strategies and ensure alignment with regulatory standards. Loser et al. [15] examined the influence of GDPR on developer priorities within agile environments, noting that security and privacy are often secondary to features that directly enhance user engagement unless mandated by law. These studies underline the significant influence of regulatory frameworks on software design and development.
@@ 2.2 Information Sources
Past research explored information sources in the field of software security [18, 19]. For example, Acar et al. [18] assessed the quality of Stack Overflow code, finding that the code was often unhelpful and insecure. Fisher et al. [19] found that 15.4% of 1.3 million analyzed Android applications contained code from Stack Overflow, with 97.9% of those containing an insecure code snippet. Similar observations have been made with privacy-relevant code [20, 21, 22]. For instance, Tahaei et al. [20] analyzed the questions developers frequently asked on Stack Overflow. They found that software developers ask for advice on privacy topics such as privacy policies, access control, and decisions on the privacy aspects of the systems they are building. In another study, Tahaei et al. [21] focused their research on the advice posted for developers regarding privacy. They found that the advice predominantly emphasizes compliance with regulations and the implementation of confidentiality measures like encryption and data minimization. Further, when analyzing developer posts on Reddit, Li et al. [22] found that developers rarely discussed privacy issues unless triggered by external factors, such as app store privacy policy updates.
@@ 2.3 Implementing Privacy Requirements
Prior work provided ample evidence that implementing complex privacy regulations is challenging or lacking [3, 23, 24, 25, 26, 27, 28]. In particular, mobile app developers repeatedly ignore privacy regulations and process personal information way beyond what they are legally allowed to [8, 9, 29, 7]. Implementing privacy measures incorrectly can threaten millions of user data, as shown by the study of Martino et al. [30] and their follow-up work in [31]. They analyzed how companies realize the data subject access request aspect of GDPR and found that due to insufficient checks of the request, it frequently can be used to leak other people’s personal information.
Balebako et al. [32] examined the privacy and security practices of smartphone app developers, particularly focusing on challenges in implementing effective privacy policies and data encryption. The research highlights a lack of awareness for privacy measures, e.g., privacy policies and data minimization, among developers, with many relying on social networks and search engines for advice.
@@ 3 Methodology
Due to the upcoming popularity of AI assistants and the lack of foundational knowledge of the underlying problem, we opted for a lab study to control the study environment fully, i.e., refraining from a second screen or device or contacting other parties with questions without us noticing. Therefore, we conducted a qualitative study with 30 professional software developers in a laboratory setting. We focused on GDPR, as it governed our German participants, and companies offering services in the EU must comply with it. While relevant privacy concepts also apply to other regulations (e.g., CCPA), the specifics of compliance may vary. Regulatory differences, such as varying data deletion rights (CCPA [2] vs. GDPR [33]) and consent requirements [34], may influence participant decisions and task performance (see Limitations 3.9 for details). Participants were divided into the following 3 groups.
@@
Expert Support (E): Third, we tested if implementation might improve with privacy expert support, addressing the communication gap found in past research [5], where developers reported a lack of support on privacy requirements. Thus, participants in the Expert-Support Group were prompted for privacy compliance and offered to contact a privacy expert through a chat function. The expert was responsible for providing individual consultations, while ensuring consistent responses to identical questions. Technical issues, such as reversing a participant’s accidental deletion of user data, were beyond the expert’s scope.
We analyzed participants’ use of privacy sources and their compliance based on the implementation’s source code. After completing their tasks, we conducted semi-structured interviews with the participants, analyzing their thoughts on privacy, privacy sources, and privacy issues they had during the implementation. While we did check if the solutions produced by the participants were functional, e.g., would run and fulfill the stated purpose of each task, the research goal was to see if privacy was considered during the task. We summarize the essential parameters of our study in Table I [35].
@@ 3.1 Study Application Prototype
Consulting one researcher with professional software development experience and another with data protection consultant experience, we designed a scenario with sensitive health data. This was rated highly relevant to real-world applications, covering a broad range of GDPR principles and requiring additional safeguards. Thus, we developed a prototype of a health application inspired by existing health apps [36, 37]. The app’s primary function was for users to schedule and manage doctor appointments. The application encompassed vital features such as a search function to find doctors by name and specialization, an appointment request system, health record management, a calendar view for appointment scheduling, a dashboard overview, and a functioning backup system. Doctors can request access to patients’ health records and, if granted, gain access to prior treatments and medication. The app already had several active users and doctors whose data was randomly generated by the tool Faker [38]. Our application dealt with sensitive user data, such as health data, which ensured that participants had to make privacy-relevant decisions. Some privacy features (e.g., consent management and automated deletion processes for the backup) were intentionally omitted to leave the implementation of privacy-sensitive tasks to the participants.
Backend Technologies: The application’s backend was developed using the Node.js [39] runtime environment, which provides a robust and scalable platform for server-side development. As our primary web application framework for Node.js, Express.js [40] facilitates the creation of robust web applications and APIs with a minimalist and flexible set of features. To optimize the deployment and management of our MySQL database and PHPMyAdmin containers [41], we used Docker Compose [42], a tool that allows the definition and running of multi-container Docker applications. As an integral part of our email processing mechanism, we used MailCatcher [43], which intercepts and displays outgoing emails from our application.
Frontend Technologies: The frontend of our application was developed using the Ionic Framework [44], a platform that enables the creation of cross-platform mobile and web technologies such as HTML, CSS, and JavaScript. As the foundational framework for our frontend application, Angular [45] was used as it provided a comprehensive and structured framework for building dynamic and interactive web applications.
@@ 3.2 Study Tasks
Participants received a task and tool description, a list of software used in the application, and an overview of the app functions and their location to make it easier for the participants to navigate the application. We included one startup task (Task~3.2) and four study tasks in the study. Task~3.2, Task~3.2, and Task~3.2 had privacy compliance issues according to the GDPR. All Tasks had clear finish conditions provided to the participants. The full text of the tasks can be found in the supplementary material (see Section 7).
@@
% % % Task 0% : Running the Application. This task guided participants through the process of starting the application. It was intended as a warm-up and to provide participants with a quick way to restart the application in case of a crash.
@@
% % % Task 1% : Repair Backup. This task showed participants what kind of data is stored by the application to ensure they were aware of the existence of the backup function relevant for Task~3.2. It required them to finish the backup functionality of the application, which did not store one of the tables. To solve this task successfully, participants had to remove the exclusion clause, which prevented one of the tables from being backed up. There were no privacy requirements for this task, so we only report the number of functional solutions.
@@
% % % Task 2% : Permanently Deactivate Account. In this task, participants were asked to process a deletion request from one of the application’s users. While removing or anonymizing all personal identifying information from the live database was sufficient to fulfill the functional aspect of the task, participants were required to delete or anonymize the data in the backup system to preserve the user’s privacy correctly.
@@
% % % Task 3% : Search Functionality. Participants were asked to finish implementing a function allowing doctors to look up past treatments for certain diseases. To fulfill the functional requirements, participants had to query the database for patients with the disease in question and display either their medical records or the medications used to treat the disease. While giving health information to a patient’s current doctor might not be a violation, the search function without any additional filters would allow all doctors in the system to receive information on every patient. A privacy-preserving solution, however, required the participants to either implement some sort of filter such that only the doctor’s own patients were displayed or to leverage the existing health record permission system. A solution where no personally identifiable information was shown, e.g., only medications given to treat a disease without the patient’s name, was also considered privacy-preserving.
@@
% % % Task 4% : Advertisements. Finally, participants were required to implement an email function that would send advertisements to the application users. To solve this task, an automatic email notification system had to be implemented, allowing the administrator to send emails to all application users. A privacy-compliant implementation requires an additional consent system, such as adding an option to the user profile to update their consent or requesting additional permissions from all users.

### v3

@@ Abstract
While protecting user data is essential, software developers often fail to fulfill privacy requirements. However, the reasons why they struggle with privacy-compliant implementation remain unclear. Is it due to a lack of knowledge, or is it because of insufficient support? To provide foundational insights in this field, we conducted a qualitative 5-hour programming study with 30 professional software developers implementing 3 privacy-sensitive programming tasks that were designed with GDPR compliance in mind. To explore if and how developers implement privacy requirements, participants were divided into 3 groups: control, privacy prompted, and privacy expert-supported. After task completion, we conducted follow-up interviews. Alarmingly, almost all participants submitted non-GDPR-compliant solutions (79/90). In particular, none of the 3 tasks were solved privacy-compliant by all 30 participants, with the non-prompted group having the lowest number of 3 out of 30 privacy-compliant solution attempts. Privacy prompting and expert support only slightly improved participants’ submissions, with 6/30 and 8/30 privacy-compliant attempts, respectively. In fact, all participants reported severe issues addressing common privacy requirements such as purpose limitation, user consent, or data minimization. Counterintuitively, although most developers exhibited minimal confidence in their solutions, they rarely sought online assistance or contacted the privacy expert, with only 4 out of 10 expert-supported participants explicitly asking for compliance confirmation. Instead, participants often relied on existing implementations and focused on implementing functionality and security first.
@@ 1 Introduction
Several regulations have been passed to reign in tech companies’ data collection practices and empower users to regain control over their personal data. The best-known ones are Japan’s Act on the Protection of Personal Information (APPI) from 2003, South Africa’s Protection of Personal Information Act (POPI Act) from 2013, the General Data Protection Regulation (GDPR) [GDPR] (2018) by the European Union, and the California Consumer Privacy Act CCPA [CCPA] (2018) by the state of California. Consequently, user privacy has become an essential requirement that software development companies must accomplish. Severe fines [SaeTheUrb+22], fear of bad publicity [CambrideAnalytica], and customer expectations [horstmann2024those] all lead to an increased focus on developing privacy-preserving software products. Still, since the introduction of e.g., the GDPR in 2018, more than 5 billion euros have been issued for violations [enforcementtracker]. While larger companies often have access to legal experts to support software developers, their communication with software development teams tasked with fulfilling these requirements is often limited, suggesting a “communication gap” [horstmann2024those]. Although smaller companies are hit harder by fines for privacy non-compliance, they are often unable to hire privacy experts due to limited budget restrictions. As a result, fulfilling privacy requirements has become a task handled solely by the software development teams.
However, past research showed that this has not (yet) translated into the actual practices of individual developers (e.g., [SaeTheUrb+22, KocAltJoh23, NguBacMarSto21, NguBacSto22, tahaei2022embedding, kekulluouglu2023we, prybylo2024evaluating]), raising the question: are developers unaware of privacy requirements, or is the task of fulfilling them lacking support? To address this research gap, we conducted a lab programming study with $30$ professional software developers implementing requirements for a health application, lasting 3 hours and 23 minutes on average. We asked participants to complete 4 programming tasks, with 3 of them requiring a GDPR-compliant implementation: (i) a warm-up task consisting of fixing an application’s backup function, (ii) processing a deletion request from one of the application’s users, (iii) implementing a search function for doctor users of the application, and lastly, (iv) implementing an e-mail function to send user advertisements and obtain user consent for a privacy-compliant solution. With this work, we explored the following research questions:
@@
To explore developers’ privacy awareness and address the communication gap [horstmann2024those], we separated the participants into three groups: (i) a control group without mentioning privacy at all, (ii) a group prompted on privacy-compliant implementation, and (iii) a chatting group having access to a privacy expert. To further understand participants’ implementation challenges with privacy requirements during the study, we conducted follow-up interviews.
With this work, we contribute foundational insights into a highly complex problem. Our findings suggest a lack of privacy awareness, the perception that privacy requirements are overly complex to implement, and a reluctance to consult privacy experts, as developers often fear being a burden or inconveniencing them. Participants cited a lack of knowledge and experience and relied on preexisting implementations. Still, most interactions with experts were limited to confirming compliance rather than seeking guidance on privacy requirements.
@@ 2.1 Regulations
Previous research [campanile2022towards, marjanov2023data, SaeTheUrb+22, loser2014security, peixoto2023perspective, alhazmi2021m] provided comprehensive insights into how GDPR and similar regulations shape the software development process. Marjanov et al. [marjanov2023data] emphasized the necessity of both technical and organizational measures to meet legal standards, identifying specific danger points and conditions that lead to data breaches. This data-driven approach offered developers crucial insights into compliance risks and mitigation strategies. Saemann et al. [SaeTheUrb+22] further explored the impact of regulations through the analysis of GDPR fines, highlighting common compliance failures, particularly during data collection. They advocated for the integration of Privacy-by-Design principles early in the development cycle to enhance data protection strategies and ensure alignment with regulatory standards. Loser et al. [loser2014security] examined the influence of GDPR on developer priorities within agile environments, noting that security and privacy are often secondary to features that directly enhance user engagement unless mandated by law. These studies underline the significant influence of regulatory frameworks on software design and development.
@@ 2.2 Information Sources
Past research explored information sources in the field of software security [acar2016you, fischer2017stack]. For example, Acar et al. [acar2016you] assessed the quality of Stack Overflow code, finding that the code was often unhelpful and insecure. Fisher et al. [fischer2017stack] found that 15.4% of 1.3 million analyzed Android applications contained code from Stack Overflow, with 97.9% of those containing an insecure code snippet. Similar observations have been made with privacy-relevant code [tahaei20understanding, tahaei2022understanding, li2021developers]. For instance, Tahaei et al. [tahaei20understanding] analyzed the questions developers frequently asked on Stack Overflow. They found that software developers ask for advice on privacy topics such as privacy policies, access control, and decisions on the privacy aspects of the systems they are building. In another study, Tahaei et al. [tahaei2022understanding] focused their research on the advice posted for developers regarding privacy. They found that the advice predominantly emphasizes compliance with regulations and the implementation of confidentiality measures like encryption and data minimization. Further, when analyzing developer posts on Reddit, Li et al. [li2021developers] found that developers rarely discussed privacy issues unless triggered by external factors, such as app store privacy policy updates.
@@ 2.3 Implementing Privacy Requirements
Prior work provided ample evidence that implementing complex privacy regulations is challenging or lacking [SaeTheUrb+22, balebako2014improving, senarath2019will, senarath2018understanding, boteju2023sok, ma2024programmer, utz2023privacy]. In particular, mobile app developers repeatedly ignore privacy regulations and process personal information way beyond what they are legally allowed to [NguBacMarSto21, NguBacSto22, KocWesAltOlv+22, KocAltJoh23]. Implementing privacy measures incorrectly can threaten millions of user data, as shown by the study of Martino et al. [MarRobWeyQua+19] and their follow-up work in [MarMeeQua+22]. They analyzed how companies realize the data subject access request aspect of GDPR and found that due to insufficient checks of the request, it frequently can be used to leak other people’s personal information.
Balebako et al. [balebako2014privacy] examined the privacy and security practices of smartphone app developers, particularly focusing on challenges in implementing effective privacy policies and data encryption. The research highlights a lack of awareness for privacy measures, e.g., privacy policies and data minimization, among developers, with many relying on social networks and search engines for advice.
@@ 3 Methodology
Due to the upcoming popularity of AI assistants and the lack of foundational knowledge of the underlying problem, we opted for a lab study to control the study environment fully, i.e., refraining from a second screen or device or contacting other parties with questions without us noticing. Therefore, we conducted a qualitative study with 30 professional software developers in a laboratory setting. We focused on GDPR, as it governed our German participants, and companies offering services in the EU must comply with it. While relevant privacy concepts also apply to other regulations (e.g., CCPA), the specifics of compliance may vary. Regulatory differences, such as varying data deletion rights (CCPA [CCPA] vs. GDPR [gdprart17]) and consent requirements [gdprccpacomp], may influence participant decisions and task performance (see Limitations 3.9 for details). Participants were divided into the following 3 groups.
@@
Expert Support (E): Third, we tested if implementation might improve with privacy expert support, addressing the communication gap found in past research [horstmann2024those], where developers reported a lack of support on privacy requirements. Thus, participants in the Expert-Support Group were prompted for privacy compliance and offered to contact a privacy expert through a chat function. The expert was responsible for providing individual consultations, while ensuring consistent responses to identical questions. Technical issues, such as reversing a participant’s accidental deletion of user data, were beyond the expert’s scope.
We analyzed participants’ use of privacy sources and their compliance based on the implementation’s source code. After completing their tasks, we conducted semi-structured interviews with the participants, analyzing their thoughts on privacy, privacy sources, and privacy issues they had during the implementation. While we did check if the solutions produced by the participants were functional, e.g., would run and fulfill the stated purpose of each task, the research goal was to see if privacy was considered during the task. We summarize the essential parameters of our study in Table I [serafini2024engaging].
@@ 3.1 Study Application Prototype
Consulting one researcher with professional software development experience and another with data protection consultant experience, we designed a scenario with sensitive health data. This was rated highly relevant to real-world applications, covering a broad range of GDPR principles and requiring additional safeguards. Thus, we developed a prototype of a health application inspired by existing health apps [zocdoc, doctolib]. The app’s primary function was for users to schedule and manage doctor appointments. The application encompassed vital features such as a search function to find doctors by name and specialization, an appointment request system, health record management, a calendar view for appointment scheduling, a dashboard overview, and a functioning backup system. Doctors can request access to patients’ health records and, if granted, gain access to prior treatments and medication. The app already had several active users and doctors whose data was randomly generated by the tool Faker [faker]. Our application dealt with sensitive user data, such as health data, which ensured that participants had to make privacy-relevant decisions. Some privacy features (e.g., consent management and automated deletion processes for the backup) were intentionally omitted to leave the implementation of privacy-sensitive tasks to the participants.
Backend Technologies: The application’s backend was developed using the Node.js [nodejs] runtime environment, which provides a robust and scalable platform for server-side development. As our primary web application framework for Node.js, Express.js [expressjs] facilitates the creation of robust web applications and APIs with a minimalist and flexible set of features. To optimize the deployment and management of our MySQL database and PHPMyAdmin containers [PHPMyAdmin], we used Docker Compose [docker], a tool that allows the definition and running of multi-container Docker applications. As an integral part of our email processing mechanism, we used MailCatcher [mailcatcher], which intercepts and displays outgoing emails from our application.
Frontend Technologies: The frontend of our application was developed using the Ionic Framework [IonicFramework], a platform that enables the creation of cross-platform mobile and web technologies such as HTML, CSS, and JavaScript. As the foundational framework for our frontend application, Angular [Angular] was used as it provided a comprehensive and structured framework for building dynamic and interactive web applications.
@@ 3.2 Study Tasks
Participants received a task and tool description, a list of software used in the application, and an overview of the app functions and their location to make it easier for the participants to navigate the application. We included one startup task (Task 3.2) and four study tasks in the study. Task 3.2, Task 3.2, and Task 3.2 had privacy compliance issues according to the GDPR. All Tasks had clear finish conditions provided to the participants. The full text of the tasks can be found in the supplementary material (see Section 7).
@@
Task 0: Running the Application. This task guided participants through the process of starting the application. It was intended as a warm-up and to provide participants with a quick way to restart the application in case of a crash.
@@
Task 1: Repair Backup. This task showed participants what kind of data is stored by the application to ensure they were aware of the existence of the backup function relevant for Task 3.2. It required them to finish the backup functionality of the application, which did not store one of the tables. To solve this task successfully, participants had to remove the exclusion clause, which prevented one of the tables from being backed up. There were no privacy requirements for this task, so we only report the number of functional solutions.
@@
Task 2: Permanently Deactivate Account. In this task, participants were asked to process a deletion request from one of the application’s users. While removing or anonymizing all personal identifying information from the live database was sufficient to fulfill the functional aspect of the task, participants were required to delete or anonymize the data in the backup system to preserve the user’s privacy correctly.
@@
Task 3: Search Functionality. Participants were asked to finish implementing a function allowing doctors to look up past treatments for certain diseases. To fulfill the functional requirements, participants had to query the database for patients with the disease in question and display either their medical records or the medications used to treat the disease. While giving health information to a patient’s current doctor might not be a violation, the search function without any additional filters would allow all doctors in the system to receive information on every patient. A privacy-preserving solution, however, required the participants to either implement some sort of filter such that only the doctor’s own patients were displayed or to leverage the existing health record permission system. A solution where no personally identifiable information was shown, e.g., only medications given to treat a disease without the patient’s name, was also considered privacy-preserving.
@@
Task 4: Advertisements. Finally, participants were required to implement an email function that would send advertisements to the application users. To solve this task, an automatic email notification system had to be implemented, allowing the administrator to send emails to all application users. A privacy-compliant implementation requires an additional consent system, such as adding an option to the user profile to update their consent or requesting additional permissions from all users.